OpenStudy (anonymous):

2sqrt{x+2}-sqrt{x-3}=sqrt{x+9}

OpenStudy (anonymous):

\[2\sqrt{x+2}-\sqrt{x-3}=\sqrt{x+9}\]

OpenStudy (anonymous):

Square all terms. It would help

OpenStudy (anonymous):

I'm still confused..

OpenStudy (anonymous):

I'm solving it... gimme a second

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[2(x+2)^.5 - (x-3)^.5 = (x+9)^.5\] \[2(x+2)^.5 = (x+9)^.5 + (x-3)^.5\] \[4(x+2) = (x+9) + (x-3) + 2(x+9)^.5(x-3)^.5\] \[4x+8=2x+6 + 2(x+9)^.5 + (x-3)^.5\] x+1= (x+9)^.5(x-3)^.5

OpenStudy (anonymous):

square both sides to get x=7 sry g2g

OpenStudy (anonymous):

well i have to go but thanks anyways